In the right triangle, find the length of the side not given a=7 b=9
Thank you

Answer by stanbon(75887)

In the right triangle, find the length of the side not given a=7 b=9
----
If c is the hypotenuse,
c^2 = 7^2+9^2
---
c^2 = 49 + 81
---
c = sqrt(130)
---
c = 11.4
Cheers,
Stan H.